Queen's Gate Inc. generates perpetual annual EBIT of $500. The cost of unlevered equity for Queen's Gate is 10%. The corporate tax rate is 35%. Queen's Gate has 300 shares outstanding that trade for $6.89. Queen's Gate has debt worth $1,512. What is the present value of financial distress costs for Queen's Gate? Assume that there are no other market imperfections other than costs of bankruptcy.
Answer #1

Queen's Gate has N = 300 shares outstanding that trade for P = $ 6.89.

Hence, Value of equity = VE = P x N = 6.89 x 300 = 2,067

Value of equity of a levered firm is also given by = Value of an unlevered firm + PV (Interest tax shield) - PV (Financial distress costs)

Value of unlevered firm = EBIT x (1 - T) / Keu = 500 x (1 - 35%) / 10% = 3,250

PV (Interest tax shield) = T x Debt = 35% x 1,512 =  529.20

Hence, VE = 2,067 = Value of an unlevered firm + PV (Interest tax shield) - PV (Financial distress costs)

= 3,250 - 529.20 - PV (Financial distress costs) =  2,720.80 - PV (Financial distress costs)

Hence, PV (Financial distress costs) = 2,720.80 - 2,067 = 653.80